OpenStudy (anonymous):

3(x + 1) = −2(x − 1) − 4

OpenStudy (anonymous):

x=-1

OpenStudy (anonymous):

Maths never sucks its just gets boring when our brain stops working * no offences meant

OpenStudy (anonymous):

Thank You So Much ^.^

OpenStudy (anonymous):

x=-1

OpenStudy (anonymous):

3x+1=-2x-1-4 3x-2x=x x=1+-1-4 x=-4

OpenStudy (anonymous):

i got x=-1

OpenStudy (anonymous):

@krista<3 there is a bracket

OpenStudy (anonymous):

im right

OpenStudy (anonymous):

@krista<3 u r wierd

OpenStudy (anonymous):

im right

OpenStudy (anonymous):

oh man u r mad y dont u apply sense anyways thats none of my concern

OpenStudy (anonymous):

I know I'm RIGHT and U R WRONG

OpenStudy (anonymous):

ohkay u r right but i am also right

OpenStudy (anonymous):

anyways talk to eminem

OpenStudy (anonymous):

no ur not and what why eminem he is a rapper

OpenStudy (anonymous):

its useless

OpenStudy (anonymous):

y eminem he is a rapper

OpenStudy (anonymous):

yes he is and he shuld talk to you

OpenStudy (anonymous):

why?

OpenStudy (anonymous):

and how?

OpenStudy (anonymous):

do you want to talk and cuz u need rest and motivation too badly

OpenStudy (anonymous):

wat?